OpenStudy (anonymous):

Which value of B solves the equation? B^-3=1/27 A. 3 B. 27 C. 9 D. 81

OpenStudy (anonymous):

I'll help

OpenStudy (anonymous):

C

OpenStudy (anonymous):

@cowgirlcrazy66, @lind3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\frac{ 1 }{ B ^{3} } = \frac{ 1 }{ 27 }\]

OpenStudy (anonymous):

B = 3

OpenStudy (anonymous):

oh sorry, not C

OpenStudy (anonymous):

I believe it's A

OpenStudy (anonymous):

Yes, it's A

OpenStudy (anonymous):

Sorrry

OpenStudy (anonymous):

If you have a negative exponent in the numerator, it will be positive in the denominator. Re-write your equation first as @faisalalif1999 showed you a few posts up.

OpenStudy (anonymous):

ok thank you a lot!!!

OpenStudy (anonymous):

If you square the 3 with -3 you will get the same exact answer with 1/27

OpenStudy (anonymous):

Yes it's A.

OpenStudy (anonymous):

thx!!!!!!

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

B=3 my friend @whitetiger2288

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

(NP)

OpenStudy (anonymous):

B = 3! Let's see how long we can keep this thread alive.

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

You have: \[B^{-3} = \frac{1}{27}\] We can re-write this as: \[\frac{1}{B^3} = \frac{1}{27}\] If we multiply both sides by 27 and B^3: \[27 \times B^3 \times \frac{1}{B^3} = \frac{1}{27}\times 27 \times B^3[\] we have \[27 = B^3\] Taking the cube-root of both sides: \[\sqrt[3]{27} = \sqrt[3]{B^3}\] we get: \[3 = B\] which is \[B = 3\]

OpenStudy (anonymous):

In english, your answer is "B is equivalent to three," or "B is equal to three."

OpenStudy (anonymous):

whitetiger what time is it were you are?

OpenStudy (anonymous):

Translated to German, that is "B ist gleich drei."

OpenStudy (anonymous):

Whitetiger I have to go to bed but I will be online all day so talk to you tomorrow!!! BYE BYE